OpenStudy (anonymous):

If I am supposed to use 2 ounces per gallon (liquids). How much would I use for a 24 oz. bottle?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

128 oz = gallon for 2 ounces 64= 1 ounce 32= 1/2 ounce 16= 1/4 ounce 8= 1/8 ounce 8 x 3=24 1/8 + 1/8+ 1/8 =3/8 ounce for 24 ounces

OpenStudy (anonymous):

would that be about a teaspoon or what?

OpenStudy (anonymous):

How does it want the answer does it ask?

OpenStudy (anonymous):

Just trying to get the correct amount. Is 3/8 of an ounce about a teaspoon?

OpenStudy (anonymous):

honestly im not sure probably google it

OpenStudy (anonymous):

THANKS!! I WILL.